Start learning →

Two dice are thrown. Chance of at least one six?

A11/36
B5/36
C13/36
D1/3
Answer & Solution
Correct answer: A. 11/36
Complement: no six has 25/36, so 1 - 25/36 = 11/36.
Practise this question with Quill — SSC CGL →

Keep practising Probability

Browse all Probability questions for SSC CGL, or explore more Quantitative Aptitude topics.

Start a Probability lesson →
Related questions